OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. -2y^2 - 8y = -6

OpenStudy (anonymous):

first step:\[-2y^2-8y+6=0,\]\[y^2+4y-3=0.\]

OpenStudy (anonymous):

Since it is asking to use the quadratic formula figure out the variables so a = -2 b = -8 and c = 6 you should have a copy of the formula to go along with this question I recommend memorizing it. First thing you should do is solve what is under the radical if you end up with a negative it there are no solutions. Remember -b = -(-8) = 8 (a negative times a negative gives a positive)